# Arithmetic of cuts in ordered abelian groups and of ideals over valuation rings

### v2

@@ Abstract.
ordered abelian group, cut, final segment, valuation ring, ideal.
@@ 1. Introduction
For the notions and notation we use in this paper, see Sections 2.1, 2.2, 3.1 and 3.2. One main purpose of this manuscript is to address a problem that comes up in the papers [3, 4]. There we work with a valued field $(K,v)$ and consider quotients $U/UV$ of ideals $U$ and $V$ of its valuation ring $\mathcal{O}_{v}\,$. We wish to compute the annihilator of the $\mathcal{O}_{v}$-module $U/UV$. In the present manuscript, we will consider the more general case of quotients $I_{1}/I_{2}$ of $\mathcal{O}_{v}$-ideals $I_{1},I_{2}$ (see Section 3.5). To this end, we consider equations $I_{1}J=I_{2}$ and inequalities $I_{1}J\subseteq I_{2}\,$, as we will describe in detail below.
@@
(2) $S_{1}\,-_{\rm fs}\,S_{2}\>:=\>\{\alpha\in G\mid\alpha+S_{1}\subseteq S_{2}\}$
is the largest solution $T_{\rm max}$ for the inequality $S_{1}+T\subseteq S_{2}\,$. However, in the literature there does not seem to exist any recipy for computing $S_{1}-_{\rm fs}S_{2}$ in terms of (elementwise) sums and differences of final segments derived from $S_{1}$ and $S_{2}$ that would help solve the problems that came up in the papers [3, 4] and which we will explain below.
A key to the answers to the above questions is the notion of invariance group of a final segment, which always is a convex subgroup of $G$ (see Section 2.4). This sheds light on the important role that convex subgroups play in nonarchimedean ordered groups for the answers to the above questions. The case of archimedean ordered groups, which have no proper nontrivial convex subgroups, is significantly easier. Our answers to the above questions are given in Theorems 2.35 and 2.36.
@@
Take a valued field $(K,v)$ with value group $vK$ and valuation ring $\mathcal{O}_{v}$. For $a\in K$ we write $va$ in place of $v(a)$. When we talk of $\mathcal{O}_{v}$-ideals we will always assume them to be nonzero, and we also include fractional ideals; in other words, we talk about nonzero $\mathcal{O}_{v}$-modules $I\subsetneq K$. Just like the set of final segments of $vK$, the set of $\mathcal{O}_{v}$-ideals is linearly ordered by inclusion, and unions and intersections of any collections of $\mathcal{O}_{v}$-ideals are again $\mathcal{O}_{v}$-ideals.. The function
@@
is the largest solution $J_{\rm max}$ for the inequality $I_{1}J\subseteq I_{2}\,$. Again, in the literature there does not seem to exist any recipy for computing $I_{1}:I_{2}$ from $I_{1}$ and $I_{2}$ that would help solve the problems addressed in Section 3.5. This section is devoted to the computation of the annihilators $I_{2}:I_{1}$ of quotients $I_{1}/I_{2}$ of $\mathcal{O}_{v}$-ideals $I_{1},I_{2}$ with $I_{2}\subsetneq I_{1}\,$. The results are then applied in Section 3.6 to compute the annihilators of some special quotients that appear in [3] and [4] and to answer the question in which cases they are equal to the maximal ideal $\mathcal{M}_{v}$ of $\mathcal{O}_{v}\,$. Our answers to the above questions are given in Theorems 3.13 and 3.14.
@@ 2.1. Preliminaries on cuts and final segments
If $\emptyset\neq M_{1}\,,\,M_{2}\subsetneq\Gamma$ are such that $M_{1}\leq M_{2}$ and $\Gamma=M_{1}\cup M_{2}$, then we call $(M_{1},M_{2})$ a quasi-cut in $\Gamma$. It follows that $M_{1}$ is an initial segment of $\Gamma$, $M_{2}$ is a final segment of $\Gamma$, and the intersection of $M_{1}$ and $M_{2}$ consists of at most one element. If this intersection is empty, then $(M_{1},M_{2})$ is a (Dedekind) cut in $\Gamma$, in which case we write $\Lambda^{L}=M_{1}\,$, $\Lambda^{R}=M_{2}\,$, and $\Lambda=(\Lambda^{L},\Lambda^{R})$.
For any subset $M\subseteq\Gamma$, we let $M^{+}$ denote the final segment
@@
That is, $M^{+}$ is the largest final segment having empty intersection with $M$. Similarly, we set
@@
We note that if $S$ is a final segment of $\Gamma$, then $S^{c}=\{\gamma\in\Gamma\mid\gamma<S\}$ is an initial segment and $(S^{c},S)$ is a cut.
@@
3) If $\,\Gamma$ is densely ordered, then an element $\gamma\in G$ is an infimum of $S$ if and only if it is a supremum of $S^{c}$.
@@
4) The final segment $S$ is closed if and only if $S$ is principal or does not have an infimum in $\Gamma$.
@@
Then $\widehat{S}$ is the closure of $S$ in $\Gamma$. We will also write $S\,\widehat{\ }\,$ for $\widehat{S}$.
@@ 2.2. Final segments in ordered abelian groups
If $S_{1}$ and $S_{2}$ are final segments, then so is $S_{1}+S_{2}\,$. With this addition, $G\uparrow$ is a commutative monoid with neutral element $0^{-}$. It is in general not a group and can even contain itempotents other than $0^{-}$; see part 1) of Lemma 2.3 and part 3) of Lemma 2.6. For example, the lexicographically ordered group $G=\mathbb{Z}\times\mathbb{Z}$ has the proper nontrivial subgroup $H=\{0\}\times\mathbb{Z}$, which gives rise to two idempotent cuts. Indeed, since $H+H=H$, the two cuts
($(\{\alpha\in G\mid\exists\beta\in H:\alpha\leq\beta\}\,,\,\{\alpha\in G\mid\forall\beta\in H:\alpha>\beta\})$
and
($(\{\alpha\in G\mid\forall\beta\in H:\alpha<\beta\}\,,\,\{\alpha\in G\mid\exists\beta\in H:\alpha\geq\beta\})$
are idempotent under both cut additions.
@@
1) Assume that $G$ is densely ordered. Then $0^{+}+0^{+}=0^{+}$.
2) Assume that $G$ is discretely ordered. If the final segment $S$ of $G$ has an infimum, then this is the smallest element of $S$ and $S$ is principal.
@@
2) If $S_{2}$ is principal but $S_{1}$ is not, then there is no final segment $T$ such that $S_{1}+T=S_{2}\,$.
3) If $S_{1}$ is a final segment of $G$ with infimum $\sigma_{1}$ and $S_{2}$ is a final segment of $G$ with infimum $\sigma_{2}\,$, then the final segment $S_{1}+S_{2}$ has infimum $\sigma_{1}+\sigma_{2}$.
@@
The proof of part 1) is straightforward. Part 2) follows immediately from part 1).
3): If both $S_{1}$ and $S_{2}$ are principal, then the assertion follows from (9). If at least one of them is nonprincipal, then by part 2) of Lemma 2.3, $G$ must be densely ordered. Now we have to consider the following two cases (up to symmetry). If $S_{1}=\sigma_{1}^{-}$ and $S_{2}=\sigma_{2}^{+}$, then by (10) and (9),
@@
If $S_{1}=\sigma_{1}^{+}$ and $S_{2}=\sigma_{2}^{+}$, then by (10) and (9) together with part 1) of Lemma 2.3,
@@
Take principal final segments $\sigma_{1}^{-}$ and $\sigma_{2}^{-}$ of $G$.
1) The function $G^{\uparrow}\ni T\mapsto\sigma_{1}^{-}+T$ is a bijection with inverse $G^{\uparrow}\ni T\mapsto(-\sigma_{1})^{-}+T$.
2) If $S$ is any final segment of $G$, then the unique final segment $T$ of $G$ such that $\sigma_{1}^{-}+T=S$ is $T=S+(-\sigma_{1})^{-}$. In particular, the unique final segment $T$ of $G$ such that $\sigma_{1}^{-}+T=\sigma_{2}^{-}$ is $T=(\sigma_{2}-\sigma_{1})^{-}$.
Analogously to what we already mentioned in the case of cuts, for certain final segments $S$, the function
@@
may not be injective. Part 1) of the previous proposition shows that the function (11) is injective if $S$ is principal. Part 3) of the following Lemma shows that if $S$ is nonprincipal, then the function (11) is never injective, provided that $G$ is nontrivial.
@@
1) We have that $0^{+}+S=\{\beta\in S\mid\alpha<\beta\mbox{ for some }\alpha\in S\}$.
2) If there is $\gamma\in G$ such that $S=\gamma^{-}$, then $0^{+}+S=\gamma^{+}$.
3) If $S$ is nonprincipal, then $0^{+}+S=S$.
4) Assume that $S_{1}$ is nonprincipal. Then $S_{1}+T=S_{1}+\widehat{T}$.
@@
The proof of statements 1) and 2) is straightforward. If $S$ is nonprincipal, then for every $\beta\in S$ there is some $\alpha\in S$ such that $\alpha<\beta$. Therefore, part 3) follows from part 1).
4): Since $S_{1}$ is nonprincipal, part 3) of Lemma 2.6 yields that $S_{1}=S_{1}+0^{+}$. If $T$ is closed in $G$, then $\widehat{T}=T$ and there is nothing to show. If $T$ has infimum $\gamma\in G\setminus T$, then $\widehat{T}=\gamma^{-}$ and $0^{+}+\widehat{T}=\gamma^{+}=T$. Thus, $S_{1}+\widehat{T}=S_{1}+0^{+}+\widehat{T}=S_{1}+T$. ∎
@@ 2.3. Subtraction of final segments
The notion of subtraction of Dedekind cuts has been defined in articles in the literature, such as [6] and [9]. When upper cut set addition is used, then for two Dedekind cuts $\Lambda_{1}=(\Lambda_{1}^{L},\Lambda_{1}^{R})$ and $\Lambda_{2}=(\Lambda_{2}^{L},\Lambda_{2}^{R})$, their difference $\Lambda_{1}-\Lambda_{2}$ has been defined to be the cut with upper cut set $\Lambda_{1}^{R}-\Lambda_{2}^{L}$, which is the sum of the final segments $\Lambda_{1}^{R}$ and $-\Lambda_{2}^{L}$ and thus again a final segment.
Analogously, we can define a difference of two final segments $S_{1}$ and $S_{2}$ of $G$ as
$S_{1}\,-\,S_{2}^{c}\>.$
This definition has for instance been used in [6, 9]. However, it does not extend the difference operation of the embedded copy of $G$. Indeed, for every $\alpha\in G$ we have $(\alpha^{-})^{c}=\{\beta\in G\mid\beta<\alpha\}$, whence
$-(\alpha^{-})^{c}\>=\>\{\gamma\in G\mid\gamma>-\alpha\}\>=\>(-\alpha)^{+}$
and therefore,
@@
In order to obtain a difference operation that extends the one of $G$, we define the (modified) final segment subtraction of two final segments $S_{1}$ and $S_{2}$ as
$S_{2}\,-_{\rm mc}\,S_{1}\>:=\>S_{2}\,-\,\{\gamma\in G\mid\gamma\leq S_{1}\}$
using the “modified complement” $\{\gamma\in G\mid\gamma\leq S_{1}\}$. This subtraction extends the subtraction of $G$: for arbitrary $\alpha,\beta\in G$,
(13) $\alpha^{-}\>-_{\rm mc}\>\beta^{-}\>=\>\alpha^{-}\>+\>(-\beta)^{-}\>=\>(\alpha-\beta)^{-}\>.$
@@
A final segment of $G$ has an additive inverse if and only if it is principal.
@@
By (13), a principal final segment $\alpha^{-}$ has additive inverse $(-\alpha)^{-}$. On the other hand, if a final segment $S$ is nonprincipal, then by part 1) of Proposition 2.4 the equation $S+T=0^{-}$ cannot have a solution $T$ since $0^{-}$ is principal. ∎
To simplify notation, for a final segment $S$ of $G$, we set
$\Delta S\>:=\>\{\gamma\in G\mid\gamma\geq-S\}\>.$
Lemma 2.8.
Take a final segment $S$ of $G$. If $S$ is principal, say $S=\alpha^{-}$ for some $\alpha\in G$, then
(14) $\Delta\alpha^{-}\>=\>(-\alpha)^{-}\>.$
If $S$ is nonprincipal, then $\Delta S=-S^{c}=\widehat{-S^{c}}$. Further, $\Delta S$ is always closed.
Proof.
Assume that $S=\alpha^{-}$ for some $\alpha\in G$. Then $-S=\{\beta\in G\mid\beta\leq-\alpha\}$ and $\Delta S=\{\gamma\in G\mid\gamma\geq-S\}=\{\gamma\in G\mid\gamma\geq-\alpha\}=(-\alpha)^{-}$.
Assume now that $S$ is nonprincipal. Then $\gamma\geq-S\Leftrightarrow-\gamma\leq S\Leftrightarrow-\gamma<S\Leftrightarrow-\gamma\in S^{c}\Leftrightarrow\gamma\in-S^{c}$, so $\Delta S=-S^{c}$. If $S=\alpha^{+}$ for some $\alpha\in G$, then $\Delta\alpha^{+}$ is closed by equation (14), so $-S^{c}=\widehat{-S^{c}}$ in this case. If $S$ has no infimum in $G$, then $S^{c}$ has no supremum and $-S^{c}$ has no infimum, hence again, $-S^{c}=\widehat{-S^{c}}$. This proves that $\Delta S$ is closed if $S$ is nonprincipal. By (14), $\Delta S$ is also closed if $S$ is principal. ∎
@@
Take three final segments $S_{1}\,$, $S_{2}$ and $S_{3}$ of $G$.
@@
2) If $S_{1}$ is nonprincipal, then
@@
3) If $S_{1}$ is principal, say $S_{1}=\alpha^{-}$ for some $\alpha\in G$, then
@@
6) We have that $\Delta(\Delta S)=\widehat{S}$.
7) The following holds:
@@
1): We compute:
$S_{2}\,-_{\rm mc}\,S_{1}\>=\>S_{2}\,+\,\{-\beta\mid\beta\leq S_{1}\}\>=\>S_{2}\,+\,\{\gamma\mid\gamma\geq-S_{1}\}\>=\>S_{2}\,+\,\Delta S_{1}\>.$
2): This follows from part 1) together with Lemma 2.8.
3): This follows from part 1) together with equation (14).
4): For $\Delta S_{1}$ this holds by definition, so for $S_{2}-_{\rm mc}S_{1}$ it follows from part 1). For $-S_{1}^{c}$ we observe that $S_{1}^{c}$ is an initial segment, hence $-S_{1}^{c}$ is a final segment and the same holds for $S_{2}-S_{1}^{c}=S_{2}+(-S_{1}^{c})$.
5): The proof is straightforward, using the equality $\alpha+(\beta-\gamma)=(\alpha-\gamma)+\beta$.
6): We compute:
$\Delta(\Delta S)$ $=$ $\Delta\{\alpha\in G\mid\alpha\geq-S\}\>=\>\{\beta\in G\mid\beta\geq-\{\alpha\in G\mid\alpha\geq-S\}\}$
$=$ $\{\beta\in G\mid\beta\geq\{\gamma\in G\mid\gamma\leq S\}\}\>=\>\widehat{S}\>.$
Here, the last equality is seen as follows. If $S$ has infimum $\sigma$ in $G$, then $\gamma\leq S\Leftrightarrow\gamma\leq\sigma$ and therefore, $\{\beta\in G\mid\beta\geq\{\gamma\in G\mid\gamma\leq S\}\}=\{\beta\in G\mid\beta\geq\sigma\}=\widehat{S}\,$. If $S$ has no infimum in $G$, then $\{\gamma\in G\mid\gamma\leq S\}=\{\gamma\in G\mid\gamma<S\}$ has no supremum in $G$, whence $\{\beta\in G\mid\beta\geq\{\gamma\in G\mid\gamma\leq S\}\}=\{\beta\in G\mid\beta>\{\gamma\in G\mid\gamma<S\}\}=S=\widehat{S}\,$.
7): Using parts 1) and 6), we compute:
$S_{2}\,-_{\rm mc}\,(0^{-}\,-_{\rm mc}\,S_{1})\>=\>S_{2}\,-_{\rm mc}\,(0^{-}+\Delta S_{1})\>=\>S_{2}\,-_{\rm mc}\,\Delta(\Delta S_{1})\>=\>S_{2}\,-_{\rm mc}\,\widehat{S}_{1}\>.$

### v3

@@ Abstract.
ordered abelian group, cut, final segment, valuation ring, ideal, quotients of ideals over valuation rings, annihilator.
@@ 1. Introduction
For the notions and notation we use in this paper, see Sections 2.1, 2.2, 3.1 and 3.2. One main purpose of this manuscript is to address a problem that comes up in the papers [3, 4]. There we work with valued field extensions $(L|K,v)$ of prime degree with unique extension of the valuation $v$ from $K$ to $L$. We show that the module of relative differentials (Kähler differentials) for the extension $\mathcal{O}_{L}|\mathcal{O}_{K}$ of the respective valuation rings is isomorphic, as $\mathcal{O}_{L}$-module, to a quotient $U/UV$ of ideals $U$ and $V$ of $\mathcal{O}_{L}\,$. We wish to compute the annihilator of such a quotient $U/UV$. It will contain $V$, but there are cases where it is bigger.
In the present manuscript, we will consider the more general case of quotients $I_{1}/I_{2}$ of $\mathcal{O}_{v}$-ideals $I_{1},I_{2}$ (see Section 3.5). To this end, we consider equations $I_{1}J=I_{2}$ and inequalities $I_{1}J\subseteq I_{2}\,$, as we will describe in detail below.
@@
(2) $S_{2}\,-_{\rm fs}\,S_{1}\>:=\>\{\alpha\in G\mid\alpha+S_{1}\subseteq S_{2}\}$
is the largest solution $T_{\rm max}$ for the inequality $S_{1}+T\subseteq S_{2}\,$; hence if a solution of (1) exists, then $S_{2}-_{\rm fs}S_{1}$ is the largest. However, in the literature there does not seem to exist any recipy for computing $S_{2}-_{\rm fs}S_{1}$ in terms of (elementwise) sums and differences of final segments derived from $S_{1}$ and $S_{2}$ that would help solve the problems that came up in the papers [3, 4] and which we will explain below.
A key to the answers to the above questions is the notion of invariance group of a final segment, which always is a convex subgroup of $G$ (see Section 2.4). This sheds light on the important role that convex subgroups play in nonarchimedean ordered groups for the answers to the above questions. The case of archimedean ordered groups, which have no proper nontrivial convex subgroups, is significantly easier. Our answers to the above questions are given in Theorems 2.33 and 2.34.
@@
Take a valued field $(K,v)$ with value group $vK$ and valuation ring $\mathcal{O}_{v}$. For $a\in K$ we write $va$ in place of $v(a)$. When we talk of $\mathcal{O}_{v}$-ideals we will always assume them to be nonzero, and we also include fractional ideals; in other words, we talk about nonzero $\mathcal{O}_{v}$-modules $I\subsetneq K$. Just like the set of final segments of $vK$, the set of $\mathcal{O}_{v}$-ideals is linearly ordered by inclusion, and unions and intersections of any collections of $\mathcal{O}_{v}$-ideals are again $\mathcal{O}_{v}$-ideals. The function
@@
is the largest solution $J_{\rm max}$ for the inequality $I_{1}J\subseteq I_{2}\,$. Again, in the literature there does not seem to exist any recipy for computing $I_{1}:I_{2}$ from $I_{1}$ and $I_{2}$ that would help solve the problems addressed in Section 3.5. This section is devoted to the computation of the annihilators $I_{2}:I_{1}$ of quotients $I_{1}/I_{2}$ of $\mathcal{O}_{v}$-ideals $I_{1},I_{2}$ with $I_{2}\subseteq I_{1}\,$. The results are then applied in Section 3.6 to compute the annihilators of some special quotients that appear in [3] and [4] and to answer the question in which cases they are equal to the maximal ideal $\mathcal{M}_{v}$ of $\mathcal{O}_{v}\,$. Our answers to the above questions are given in Theorems 3.13 and 3.14.
@@ 2.1. Preliminaries on cuts and final segments
If $\emptyset\neq M_{1}\,,\,M_{2}\subsetneq\Gamma$ are such that $M_{1}\leq M_{2}$ and $\Gamma=M_{1}\cup M_{2}$, then we call $(M_{1},M_{2})$ a quasi-cut in $\Gamma$, and we write $\Lambda^{L}=M_{1}\,$, $\Lambda^{R}=M_{2}\,$, and $\Lambda=(\Lambda^{L},\Lambda^{R})$. It follows that $M_{1}$ is an initial segment of $\Gamma$, $M_{2}$ is a final segment of $\Gamma$, and the intersection of $M_{1}$ and $M_{2}$ consists of at most one element. If this intersection is empty, then $(M_{1},M_{2})$ is called a (Dedekind) cut in $\Gamma$.
For any subset nonempty $M\subseteq\Gamma$, we set
@@
If this is nonempty, it is the largest final segment having empty intersection with $M$. Similarly, we set
@@
We note that if $S\subsetneq\Gamma$ is a final segment of $\Gamma$, then $S^{c}=\{\gamma\in\Gamma\mid\gamma<S\}$ is an initial segment and $(S^{c},S)$ is a cut.
@@
3) If $\,\Gamma$ is densely ordered, then an element $\gamma\in G$ is an infimum of $S$ if and only if it is a supremum of $S^{c}$. ∎
@@
4) The final segment $S$ is closed if and only if $S$ is principal or does not have an infimum in $\Gamma$. ∎
@@
Then $\widehat{S}$ is the closure of $S$ in $\Gamma$. We will also write $S\,\widehat{\ }\,$ for $\widehat{S}$ in case the expression $S$ is too long to accommodate a hat.
@@ 2.2. Final segments in ordered abelian groups
If $S_{1}$ and $S_{2}$ are final segments, then so is $S_{1}+S_{2}\,$. With this addition, $G\uparrow$ is a commutative monoid with neutral element $0^{-}$. It is in general not a group and can even contain itempotents other than $0^{-}$; see part 1) of Lemma 2.3 and part 3) of Lemma 2.6. For example, the lexicographically ordered group $G=\mathbb{Z}\times\mathbb{Z}$ has the proper nontrivial subgroup $H=\{0\}\times\mathbb{Z}$, which gives rise to two idempotent final segments. Indeed, since $H+H=H$, both
$\{\alpha\in G\mid\forall\beta\in H:\alpha>\beta\})\quad\mbox{ and }\quad\{\alpha\in G\mid\exists\beta\in H:\alpha\geq\beta\})$
are idempotent, and it can be shown that together with $G$ itself they are the only idempotent final segments in $G$.
@@
1) We have that $G$ is densely ordered if and only if $0^{+}+0^{+}=0^{+}$.
2) Assume that $G$ is discretely ordered. If the final segment $S$ of $G$ has an infimum, then this is the smallest element of $S$ and $S$ is principal. ∎
@@
2) If $S_{2}$ is principal but $S_{1}$ is not, then equation (1) has no solution.
3) If $S_{1}$ has infimum $\alpha_{1}$ and $S_{2}$ has infimum $\alpha_{2}\,$, then $S_{1}+S_{2}$ has infimum $\alpha_{1}+\alpha_{2}$.
4) If $S$ is a final segment of $G$ and $\alpha\in G$, then
(11) $\alpha+S\>=\>\alpha^{-}+S\>.$
@@
The proof of parts 1) and 4) are straightforward. Part 2) follows immediately from part 1).
3): If both $S_{1}$ and $S_{2}$ are principal, then the assertion follows from (9). If at least one of them is nonprincipal, then by part 2) of Lemma 2.3, $G$ must be densely ordered. Now we have to consider the following two cases (up to symmetry). If $S_{1}=\alpha_{1}^{-}$ and $S_{2}=\alpha_{2}^{+}$, then by (10) and (9),
@@
If $S_{1}=\alpha_{1}^{+}$ and $S_{2}=\alpha_{2}^{+}$, then by (10) and (9) together with part 1) of Lemma 2.3,
@@
Take elements $\alpha$, $\alpha_{1}^{-}$ and $\alpha_{2}^{-}$ in $G$.
1) The function
(12) $G^{\uparrow}\ni T\>\mapsto\>\alpha^{-}+T\in G^{\uparrow}$
is a bijection with inverse $G^{\uparrow}\ni T\mapsto(-\alpha)^{-}+T\in G^{\uparrow}$.
2) If $S$ is any final segment of $G$, then the unique final segment $T$ of $G$ such that $\alpha^{-}+T=S$ is $T=S+(-\alpha)^{-}=S-\alpha$. Consequently, if $S_{1}$ and $S_{2}$ are final segments of $G$ and $S_{1}$ is principal, say $S_{1}=\alpha^{-}$, then equation (1) has the unique solution
(13) $T\>=\>S_{2}\,-_{\rm fs}\,S_{1}\>=\>S_{2}-\alpha\>.$
Part 1) of this proposition shows that the function
@@
is injective if $S\in G^{\uparrow}$ is principal. Part 3) of the following Lemma shows that if $S$ is nonprincipal, then the function (14) is never injective, provided that $G$ is nontrivial.
@@
1) We have that $0^{+}+T=\{\beta\in T\mid\alpha<\beta\mbox{ for some }\alpha\in T\}$.
2) If there is $\alpha\in G$ such that $T=\alpha^{-}$, then $0^{+}+T=\alpha^{+}$.
3) If $T$ is nonprincipal, then $0^{+}+T=T$.
4) If $S$ is nonprincipal, then $S+T=S+\widehat{T}$.
@@
The proof of statements 1) and 2) is straightforward. If $T$ is nonprincipal, then for every $\beta\in T$ there is some $\alpha\in T$ such that $\alpha<\beta$. Therefore, part 3) follows from part 1).
4): Since $S$ is nonprincipal, part 3) yields that $S=S+0^{+}$. If $T$ is closed in $G$, then $\widehat{T}=T$ and there is nothing to show. If $T$ is closed in $G$, then by part 4) of Lemma 2.2, $T$ has an infimum $\alpha\in G\setminus T$. Thus, $\widehat{T}=\alpha^{-}$ and $0^{+}+\widehat{T}=\alpha^{+}=T$, so that $S+\widehat{T}=S+0^{+}+\widehat{T}=S+T$. ∎
Corollary 2.7.
For every $\alpha\in G$ and nonprincipal final segment $T$ of $G$,
(15) $\alpha+T\>=\>\alpha^{-}+T\>=\>\alpha^{+}+T\>.$
Hence the function
(16) $T\>\mapsto\>\alpha^{+}+T$
is a bijection on the set of all nonprincipal final segments of $G$, and there it is equal to the function (12).
Proof.
The first equality of (15) follows from equation (11). The second equality holds since $\alpha^{-}+T=\alpha^{-}+(0^{+}+T)=(\alpha^{-}+0^{+})+T)=\alpha^{+}+T$ by part 3) of Lemma 2.6 and equation (10). The final statement follows from the second equality of (15) together with part 1) of Proposition 2.5 and part 1) of Proposition 2.4. ∎
@@ 2.3. Subtraction of final segments
The notion of subtraction of Dedekind cuts has been defined in articles in the literature, such as [6] and [9]. When upper cut set addition is used, then for two Dedekind cuts $\Lambda_{1}=(\Lambda_{1}^{L},\Lambda_{1}^{R})$ and $\Lambda_{2}=(\Lambda_{2}^{L},\Lambda_{2}^{R})$, their difference $\Lambda_{2}-\Lambda_{1}$ has been defined to be the cut with upper cut set $\Lambda_{2}^{R}-\Lambda_{1}^{L}$, which is the sum of the final segments $\Lambda_{2}^{R}$ and $-\Lambda_{1}^{L}$ and thus again a final segment.
Analogously, we can define a difference of two final segments $S_{2}$ and $S_{1}$ of $G$ as
(18) $S_{2}\,-\,S_{1}^{c}\>=\>S_{2}\,+\,(-S_{1}^{c})\>=\>S_{2}\,+\,\{\gamma\in G\mid\gamma>-S_{1}\}\>.$
Definition 18) has for instance been used in [6, 9]. However, it does not extend the difference operation of the embedded copy of $G$. Indeed, for every $\alpha\in G$ we have
@@
In order to obtain a difference operation that extends the one of $G$, replacing “$>$” by $\geq$” in (18) we define the (modified) final segment subtraction of two final segments $S_{2}$ and $S_{1}$ as
$S_{2}\,-_{\rm mc}\,S_{1}\>:=\>S_{2}\,+\,\{\gamma\in G\mid\gamma\geq-S_{1}\}\>.$
In other words, we are replacing the complement $S_{1}^{c}$ in (18) by the “modified complement” $\{\gamma^{\prime}\in G\mid\gamma^{\prime}\leq S_{1}\}$. This subtraction extends the subtraction of $G$: for arbitrary $\alpha,\beta\in G$,
(19) $\alpha^{-}\,-_{\rm mc}\,\beta^{-}\,=\,\alpha^{-}\,+\,\{\gamma\mid\gamma\geq-\beta^{-}\}\,=\,\alpha^{-}\,+\,\{\gamma\mid\gamma\geq-\beta\}\,=\,(\alpha-\beta)^{-}\,.$
To simplify notation, for a final segment $S$ of $G$ we set
$\Delta S\>:=\>\{\gamma\in G\mid\gamma\geq-S\}\>=\>\widehat{-S^{c}}\>.$
@@
1) For each $\alpha\in G$,
(20) $\Delta\alpha^{-}\>=\>(-\alpha)^{-}\>.$
2) If $S$ is a nonprincipal final segment of $G$, then $\Delta S=-S^{c}$.
3) For every final segment $S$ of $G$, we have $\Delta(\Delta S)=\widehat{S}$.
@@
1): We have $\Delta\alpha^{-}=\{\gamma\in G\mid\gamma\geq-\alpha^{-}\}=\{\gamma\in G\mid\gamma\geq-\alpha\}=(-\alpha)^{-}$.
2): If $S$ is a nonprincipal final segment of $G$, then $\gamma\geq-S\Leftrightarrow-\gamma\leq S\Leftrightarrow-\gamma<S\Leftrightarrow-\gamma\in S^{c}\Leftrightarrow\gamma\in-S^{c}$, so $\Delta S=-S^{c}$.
3): If $S=\alpha^{-}$ for some $\alpha\in G$, then $\Delta(\Delta S)=\Delta((-\alpha)^{-})=\alpha^{-}=S=\widehat{S}$, where we have used part 1) twice. If $S$ is a nonprincipal final segment of $G$, then $\Delta(\Delta S)=\Delta(-S^{c})=(-(-S^{c})^{c})\,\widehat{\ }\,=(-(-S))\,\widehat{\ }\,=\widehat{S}$, where we have used part 2) and the definition of $\Delta$. ∎
@@
Take final segments $S_{1}\,$ and $S_{2}$ of $G$.
@@
2) The sets $\Delta S_{1}$, $S_{2}-_{\rm mc}S_{1}\,$, $-S_{1}^{c}$ and $S_{2}-S_{1}^{c}$ are again final segments of $G$.
3) If $S_{1}$ is nonprincipal, then
@@
4) If $S_{1}$ is principal, say $S_{1}=\alpha^{-}$ for some $\alpha\in G$, then
@@
6) The following holds:
@@
1): This follows from the definitions of $S_{2}-_{\rm mc}S_{1}$ and $\Delta S_{1}\,$.
2): For $\Delta S_{1}$ this holds by definition, so for $S_{2}-_{\rm mc}S_{1}$ it follows from part 1). For $-S_{1}^{c}$ we observe that $S_{1}^{c}$ is an initial segment, hence $-S_{1}^{c}$ is a final segment and the same holds for $S_{2}-S_{1}^{c}=S_{2}+(-S_{1}^{c})$.
3): This follows from part 1) together with part 2) of Lemma 2.8.
4): This follows from part 1) together with equation (20) and part 4) of Proposition 2.4.
5): The proof is straightforward, using part 1).
6): Using parts 1) and part 3) of Lemma 2.8, we compute:
$S_{2}\,-_{\rm mc}\,(0^{-}\,-_{\rm mc}\,S_{1})\>=\>S_{2}\,-_{\rm mc}\,\Delta S_{1}\>=\>S_{2}\,+\,\Delta(\Delta S_{1})\>=\>S_{2}\,+\,\widehat{S}_{1}\>.$